# Black Mirrors: CPT-Symmetric Alternatives to Black Holes

### v1

@@ A.3 In Eddington-Finkelstein coordinates
Now, as before, we can switch from $r\to\sigma$ by making the substitution (16) in Eqs. (25, 26, 26f, 26g) and extending $\sigma$ to negative values. If we now let the coordinates $v_{+}$ and $\sigma$ (respectively, $v-$ and $\sigma$) extend from $\infty$- ∞ to +$\infty$+ ∞, we obtain the general stationary black mirror, described in ingoing coordinates $(v_{+},\sigma,\theta,\hat{\varphi})$ (respectively, outgoing coordinates $(v_{-},\sigma,\theta,\hat{\varphi})$) which cover the entire spacetime except for the black horizon (respectively, white horizon). Taken together, these ingoing and outgoing coordinate systems provide an atlas covering the complete stationary black mirror spacetime.
@@
Note that, in these $(v_{\pm},\sigma,\theta,\hat{\varphi})$ coordinates, the components of the resulting metric $g_{\mu\nu}$ and field strength $F_{\mu\nu}$, as well as the components of all curvature invariants (such as the Kretschmann scalar $R_{\alpha\beta\gamma\delta}R^{\alpha\beta\gamma\delta}$) are again everywhere explicitly smooth, analytic and finite, so the spacetime has no curvature singularities. But, as in the Schwarzschild case, two of the eigenvalues $\lambda_{\pm}$ of the metric $g_{\mu\nu}$ are exchanged as we cross the horizon, $\lambda_{+}(\sigma)=-\lambda_{-}(-\sigma)$, and both of these eigenvalues have simple analytic zeros at $\sigma=0$, while the corresponding eigenvalues $\lambda_{\pm}^{-1}$ of $g^{\mu\nu}$ have simple poles).
Appendix B Saddle points of the action
@@
[…]
@@
$ds^{2}=-a^{2}(x)\,dt^{2}+dx^{2}+r^{2}(x)d\Omega^{2}$ (34)
with
${r^{\prime}}^{2}$ $=$ $1-\frac{2m}{r}+\frac{e^{2}}{r^{2}}-\Lambda r^{2},$ (35a)
$a$ $=$ $r^{\prime}.$ (35b)
The function $a=a(r)$ has at most four roots: $r_{-}$ which corresponds to the Cauchy horizon, $r_{+}$ which corresponds to the event horizon and $r_{c}$ which corresponds to the cosmological horizon, with $r_{c}=+\infty$ if $\Lambda\leq 0$. Since $a^{2}\geq 0$ it follows that the range of $r$ lies in the interval $[r_{+},r_{c}]$. Moreover, CPT symmetry implies that $r$ obeys $r^{\prime}(0)=0$, and therefore $r(0)$ is a root of $a=a(r)$. If $r(0)=r_{c}$ then the function $r=r(x)$ is a constant. Thus, the only physical initial conditions for $r$ are $r(0)=r_{+}$ and $r^{\prime}(0)=0$. Extending this solution to the horizon by changing coordinates from Schwarzschild time $t$ to Eddington-Finkelstein time $v$ ($v_{+}$ or $v_{-}$) once again brings us back to the black mirror geometry discussed in the main text.
Appendix C CPT symmetry
In the body of the paper, we focused for simplicity on the Schwarzschild (static, uncharged) black mirror, but let us now consider the general stationary black mirror ${\cal M}$ (including charge), to understand the action of CPT symmetry more clearly. For simplicity, assume the mirror has electric charge $q_{e}$ but no magnetic charge.
The natural CPT transformation Sewell (1982) maps a point with coordinates $(t,x,\theta,\varphi)\to(t,-x,\pi-\theta,\varphi+\pi)$ or, equivalently, $(t,\sigma,\theta,\varphi)\to(t,-\sigma,\pi-\theta,\varphi+\pi)$. This is an isometry that leaves both the line element $ds^{2}=g_{\mu\nu}dx^{\mu}dx^{\nu}$, as well as the field strength $F=F_{\alpha\beta}dx^{\alpha}\wedge dx^{\beta}$ invariant. However, due to the inversion of the orientation of the two-sphere, the net charge calculated via Gauss’s law from the net flux through this two sphere changes sign under this transformation. In this sense, it not only inverts spacetime, but also reverses electric charges, and is thus a CPT transformation.
This isometry maps e.g. a point $P$ with angular coordinate $(\theta,\varphi)$ near the upper corner of exterior $I$ to a mirror point $P^{\prime}$ near with angular coordinate $(\pi-\theta,\varphi+\pi)$ near the lower corner of exterior $I^{\prime}$ (see Fig. 5).
The trajectory $x^{\mu}(\tau)$ of particle of charge $q$ in the black mirror spacetime, satisfies the equation of motion
$\ddot{x}^{\mu}+\Gamma^{\mu}_{\alpha\beta}\dot{x}^{\alpha}\dot{x}^{\beta}=qF^{\mu}_{\;\;\,\nu}\,\dot{x}^{\nu}$ (36)
where a dot denotes a proper time derivative $d/d\tau$. Performing a $CPT$ transformation on the background also maps the trajectory $x^{\mu}(\tau)$ to a new trajectory $x^{\prime}{}^{\mu}(\tau)$, see Fig. 5. Moreover, since this map reverses the orientation on the $(\theta,\varphi)$ 2-spheres, it not only reverses the charge of the black hole (as measured by Gauss’s law using a sphere surrounding the hole), it also reverses the charge of the point particle (again as measured by Gauss’s law using a sphere surrounding the point charge). This is why, in Figs. 3,4, and 5, the arrows showing charge flow (more correctly, the direction of $q\dot{x}^{\mu}$) in the left-hand (red) trajectories are reversed relative to the corresponding arrows on the right-hand (blue) trajectories. [^3] This insight, that an antiparticle is just a particle moving backward in time – i.e. a particle whose “worldline time” (proper time) has been reversed relative to the ”embedding time” (the time direction in the target space) is due to Stueckelberg Stueckelberg (1941a, b, c).

### v2

@@ A.3 In Eddington-Finkelstein coordinates
Now, as before, we can switch from $r\to\sigma$ by making the substitution (14) in Eqs. (23, 24, 24f, 24g) and extending $\sigma$ to negative values. If we now let the coordinates $v_{+}$ and $\sigma$ (respectively, $v_{-}$ and $\sigma$) extend from $\infty$- ∞ to +$\infty$+ ∞, we obtain the general stationary black mirror, described in ingoing coordinates $(v_{+},\sigma,\theta,\hat{\varphi})$ (respectively, outgoing coordinates $(v_{-},\sigma,\theta,\hat{\varphi})$) which cover the entire spacetime except for the black horizon (respectively, white horizon). Taken together, these ingoing and outgoing coordinate systems provide an atlas covering the complete stationary black mirror spacetime.
@@
Note that, in these $(v_{\pm},\sigma,\theta,\hat{\varphi})$ coordinates, the components of the resulting metric $g_{\mu\nu}$ and field strength $F_{\mu\nu}$, as well as the components of all curvature invariants (such as the Kretschmann scalar $R_{\alpha\beta\gamma\delta}R^{\alpha\beta\gamma\delta}$) are again everywhere explicitly smooth, analytic and finite, so the spacetime has no curvature singularities. But, as in the Schwarzschild case, two of the eigenvalues $\lambda_{\pm}$ of the metric $g_{\mu\nu}$ are exchanged as we cross the horizon, $\lambda_{+}(\sigma)=\lambda_{-}(-\sigma)$, and both of these eigenvalues have simple analytic zeros at $\sigma=0$, while the corresponding eigenvalues $\lambda_{\pm}^{-1}$ of $g^{\mu\nu}$ have simple poles).
Appendix B CPT symmetry
In the body of the paper, we focused for simplicity on the uncharged (Schwarzschild) black mirror but, to understand the action of $CPT$ more clearly, let us now consider the charged (Reissner-Nordstrom) black mirror. The natural $CPT$ transformation of the black hole background (the unique involutive isometry which commutes with all of the continuous isometries, and has no fixed points) Gibbons (1986) sends
$(t,\sigma,\theta,\varphi)\to(t,-\sigma,\pi-\theta,\pi+\varphi).$ (25)
This isometry maps e.g. a point $P$ near the upper corner of exterior $I$ (and with the angular coordinates $\theta,\varphi$) to a mirror point $P^{\prime}$ near the lower corner of exterior $I^{\prime}$ (with antipodal angular coordinates $\pi-\theta,\pi+\varphi$); see Fig. 5.
The black mirror is obtained from the corresponding black hole by identifying each point on the BH (WH) horizon bounding exterior $I$ with its CPT image on the BH (WH) horizon bounding exterior $I^{\prime}$.
B.1 Orientations and $CPT$ versus $PT$
In this subsection, let us explain why the isometry $\varphi_{CPT}:(t,\sigma,\theta,\varphi)\to(t,-\sigma,\pi-\theta,\pi+\varphi)$ (which inverts the angular 2-sphere) is a classical analogue of the $CPT$ transformation, whereas the isometry $\varphi_{PT}:(t,\sigma,\theta,\varphi)\to(t,-\sigma,\theta,\varphi)$ (which does not invert the angular 2-sphere) corresponds to $PT$.
Figure 7: The time orientation in (a) a black hole and (b) a black mirror.
We start by reviewing the space, time, and spacetime orientations of the maximally-extended Schwarzschild black hole. In null Kruskal coordinates $(U,V,\theta,\varphi)$, the line element is
$ds^{2}=-\frac{32m^{3}}{r}{\rm e}^{-r/2m}dUdV+r^{2}d\Omega^{2}$ (26)
with $r=r(U,V)$ given by
$r=2m\left[1+W_{0}\left(-\frac{UV}{e}\right)\right]$ (27)
where $W_{0}$ is the Lambert $W$ function, and $U$ and $V$ both run from $\infty$- ∞ to +$\infty$+ ∞. This spacetime has a global 4-orientation, i.e., an everywhere non-vanishing 4-form $\mu_{4}$ (the spacetime volume 4-form). In terms of the above coordinates it is given by
$\mu_{4}=\frac{16m^{3}}{r}{\rm e}^{-r/2m}r^{2}{\rm sin}\,\theta\,dU\wedge dV\wedge d\theta\wedge d\varphi.$ (28)
To determine whether an ordered tetrad of vectors is positively or negatively oriented, we contract it with $\mu_{4}$ and check the sign of the result. In this way, we see that the tetrad of Kruskal coordinate vectors $(\partial_{U},\partial_{V},\partial_{\theta},\partial_{\varphi})$ is everywhere positively oriented (except at $\theta=0$ or $\pi$, where the spherical coordinates $\theta,\varphi$ on the 2-sphere have their usual coordinate singularity); and we can also check that the tetrad $(\partial_{t},\partial_{r},\partial_{\theta},\partial_{\varphi})$ is positively oriented in both exteriors, $I$ and $I^{\prime}$.
The maximally-extended Schwarschild black hole is foliated by global Cauchy surfaces and has a global time orientation – i.e. an everywhere non-vanishing timelike vector field $T$ (which may be taken to consist of the timelike unit normal vectors to the previously-mentioned Cauchy surfaces). For $r\gg 2m$, $T$ agrees with $+\partial_{t}$ in exterior $I$, and $-\partial_{t}$ in exterior $I^{\prime}$ (see Fig. 7a). The three-form $\mu_{3}$ obtained by contracting the time-orientation vector $T$ with the 4-orientation $\mu_{4}$ is then the 3-orientation on the black hole Cauchy surfaces.
Under the $PT$ isometry $\varphi_{PT}$, one can check that: (i) the push-forward of the vector field $T$ is $-T$; (i) the pullback of the spatial orientation $\mu_{3}$ is $-\mu_{3}$; and (iii) the pullback of the spacetime orientation $\mu_{4}$ is $+\mu_{4}$. In other words, $\varphi_{PT}$ inverts the time orientation and the spatial orientation, but preserves the spacetime orientation (as we would expect for $PT$).
Now, let us return to the more general (charged) black hole, and its electromagnetic field strength two-form $F$. Using Gauss’s law, we obtain the electric charge of the black hole (in one exterior) by integrating $\star F$ over an angular 2-sphere (in that exterior). Under the $PT$ isometry $\varphi_{PT}$, the pullback of $\star F$ is $\star F$, whereas, under the $CPT$ isometry $\varphi_{CPT}$, the pullback of $\star F$ is $-(\star F)$. So $\varphi_{PT}$ preserves the hole’s charge, while $\varphi_{CPT}$ reverses it (again, as we would expect for $PT$ and $CPT$).
B.2 Total charge (and mass) of the black hole.
In the following subsection – Appendix B.3 – we will show that the two sides of the black mirror have the same electric charge and mass, as measured by Gauss’s Law.
But first, in the present subsection, we review the corresponding result for the black hole: the two sides of the black hole have opposite electric charge (or mass), as measured by Gauss’s Law, so that the total charge (or mass) of the maximally-extended (two-sided) black hole (again, as measured by Gauss’s Law, and including the contributions from both exteriors) is zero.
One way to obtain the black hole’s total electric charge is to integrate the electric charge density over a complete Cauchy surface, e.g. the spatial 3-surface $\Sigma$ shown in Fig. 7a. But we are dealing with a vacuum solution where the charge density vanishes everywhere on this Cauchy surface, so the total electric charge must vanish as well!
We can obtain the same conclusion via Gauss’s law. In this approach, the charge is obtained by integrating $\star F$ over the Cauchy surface’s spatial 2-boundary $\partial\Sigma$. This boundary has two components: $\partial\Sigma_{+}$ (the $S^{2}$ at $r=\infty$ in region $I$) and $\partial\Sigma_{-}$ (the $S^{2}$ at $r=\infty$ in region $I^{\prime}$). When expressed in $(t,r,\theta,\varphi)$ coordinates, we see that $\star F\propto q\,{\rm sin}\,\theta\,d\theta\wedge d\varphi$ is identical in the two exteriors; but the two boundary components $\partial\Sigma_{+}$ and $\partial\Sigma_{-}$ have opposite orientations [^3] To see that $\partial\Sigma_{+}$ and $\partial\Sigma_{-}$ have opposite orientations, note that, when expressed in terms of $(t,r,\theta,\varphi)$ coordinates, the 4-orientation $\mu_{4}=r^{2}{\rm sin}\,\theta\,dt\wedge dr\wedge d\theta\wedge d\varphi$ is identical in the two exteriors; and the orientation two-form $\mu_{2}$ on the 2-boundary $\partial\Sigma$ is obtained by contracting this 4-orientation $\mu_{4}$ with the outward pointing spatial normal vector $\partial_{r}$ (the same in both exteriors), and the time orientation vector $T$, which is $+\partial_{t}$ in exterior $I$ and $-\partial_{t}$ in exterior $I^{\prime}$; so the two orientation $\mu_{2}$ is $\propto{\rm sin}\,\theta\,d\theta\wedge d\varphi$ in exterior $I$, and $\propto-{\rm sin}\,\theta\,d\theta\wedge d\varphi$ in exterior $I^{\prime}$. so the two integrals cancel.
A closely analogous argument shows that the total Komar mass of the maximally-extended (two-sided) black hole is zero. Again, in this case, the argument can be expressed either in terms of integrating the mass density over the complete 3D Cauchy surface $\Sigma$, or by using Gauss’s Law/Stokes’ Theorem to rephrase this in terms of the Cauchy surface’s two-component 2D boundary $\partial\Sigma=\partial\Sigma_{+}+\partial\Sigma_{-}$. Consider the $2D$ ($\partial\Sigma$) version of the argument: just as we computed the total (2-sided) electric charge $Q_{tot}$ by starting with the potential 1-form $A=-(q_{e}/r)dt$, computing its field strength $F_{A}=dA$, and finding $Q_{tot}\propto\int_{\partial\Sigma_{+}}\star F_{A}+\int_{\partial\Sigma_{-}}\star F_{A}=0$, we can compute the total (2-sided) Komar mass $M_{tot}$ by starting with the Killing one-form $k=-(1-\frac{2m}{r})dt$ (the one-form dual to the timelike Killing vector $\partial_{t}$), computing its “field strength” $F_{k}=dk$, and finding $M_{tot}\propto\int_{\partial\Sigma_{+}}\star F_{k}+\int_{\partial\Sigma_{-}}\star F_{k}=0$.
So, in this sense (with respect to its global time and spacetime orientation) the total charge (or mass) of the black hole vanishes; or, stated another way, its two exteriors have equal but opposite charge (or mass).
B.3 Total charge (and mass) of the black mirror.
Finally, in this section, we turn to the black mirror and find that something surprising happens, due to its non-trivial topology: as determined via Gauss’s Law from the electric field (or gravitational field) flux through a distant two-sphere, the two sides now have the same charge (or mass), and hence a total two-sided charge (or mass) which does not cancel to zero (as it did in the two-sided black hole), but is rather twice that of an ordinary (one-sided) black hole; and this is true even though, contrary to Gauss’s Law, the solution is a smooth (but topologically non-trivial) vacuum solution, with no “enclosed” electric charge (or mass) localized on the mirror. This is a novel incarnation of Misner and Wheeler’s idea of “mass without mass” and “charge without charge” Misner and Wheeler (1957); Sorkin (1977); Giulini (2018).
To understand this, let us denote the black mirror’s spacetime 4-orientation by $\tilde{\mu}_{4}$ and its electromagnetic field strength by $\tilde{F}_{A}$. Note that, in the coordinates $\{v_{\pm},\sigma,\theta,\varphi\}$ or $\{t,\sigma,\theta,\varphi\}$ in which the mirror’s two-sided metric $\tilde{g}_{\mu\nu}$ is smooth, its spacetime 4-orientation $\tilde{\mu}_{4}$ and electromagnetic field strength $\tilde{F}$ are also smooth: they are given by $\tilde{\mu}_{4}=2r^{2}\sigma\,{\rm sin}\,\theta\,dt\wedge d\sigma\wedge d\theta\wedge d\varphi$ and $\tilde{F}_{A}=-2(q\sigma/r^{2})dt\wedge d\sigma$. In fact, if we express the black hole quantities $\mu_{4}$, $F_{A}$ and $\star F_{A}$ in these same coordinates, we see that they agree with their black mirror counterparts $\tilde{\mu}_{4}$, $\tilde{F}_{A}$ and $\star{\tilde{F}}_{A}$ in the exterior regions $I$ and $I^{\prime}$. But note that the black mirror quantities $\tilde{\mu}_{4}$, $\tilde{F}_{A}$ and $\star{\tilde{F}}_{A}$ all vanish at the mirror itself (i.e at $\sigma=0$), a sign that something topologically non-trivial happens at the mirror, as we already knew from Fig. 2 and the metric eigenvalues $\lambda_{\pm}$ (6) smoothly swapping signs at $\sigma=0$.
Next note that the black mirror and black hole have different global time orientations. This may be understood from Fig. 3: from the standpoint of the black hole’s global time orientation $T$ (see Fig. 7a), point $P$ is to the future of point $Q$, whereas point $P^{\prime}$ is to the past of point $Q^{\prime}$. But, since we identify $P\sim P^{\prime}$ and $Q\sim Q^{\prime}$ in the black mirror, in order for them to have a consistent time ordering, we learn that the black mirror must have a different global time orientation $\widetilde{T}$ (see Fig. 7b). In particular, the black mirror’s time orientation $\widetilde{T}$ agrees with the black hole’s time orientation $T$ in exterior $I$, but differs from $T$ by a sign in exterior $I^{\prime}$.
Now let us consider the 2-orientation on the black mirror’s angular 2-boundary $\partial\Sigma=\partial\Sigma_{+}+\partial\Sigma_{-}$. In the previous subsection, we computed the black hole’s corresponding 2-orientation $\mu_{2}$ by contracting its spacetime 4-orientation $\mu_{4}$ with its time orientation $T$ and the outward pointing normal $\partial_{r}$. Analogously, we now compute the black mirror’s 2-orientation $\tilde{\mu}_{2}$ by contracting its spacetime 4-orientation $\tilde{\mu}_{4}$ with its time orientation $\tilde{T}$ and the outward pointing normal $\partial_{r}$. Since $T$ and $\tilde{T}$ differ by a sign in exterior $I^{\prime}$, $\mu_{2}$ and $\tilde{\mu}_{2}$ also differ by a sign in in exterior $I^{\prime}$. In other words, in the black mirror, $\star\tilde{F}$ is again $\propto+{\rm sin}\,\theta\,d\theta\wedge d\varphi$ on both sides, but now (in contrast to the black hole case) the mirror’s angular 2-orientation is also $\propto+{\rm sin}\,\theta\,d\theta\wedge d\varphi$ on both sides.
So if we use Gauss’s law to determine the mirror’s total two-sided charge from the measured electric field flux, $\tilde{Q}_{tot}\propto\int_{\partial\Sigma_{+}}\star\tilde{F}_{A}+\int_{\partial\Sigma_{-}}\star\tilde{F}_{A}$, we find that the two contributions are equal rather than opposite; and hence, instead of the total 2-sided charge canceling (as in the black hole case) it is non-zero, and twice the charge of a one-sided black hole: $\tilde{Q}_{tot}\propto 2\int_{\partial\Sigma_{+}}\star\tilde{F}_{A}=2Q_{{\rm 1-sided}}$.
In other words, since the black mirror’s time orientation is flipped in exterior $I^{\prime}$ (relative to the black hole), its charge is also flipped in exterior $I^{\prime}$ (relative to the black hole). This is consistent with the point, first emphasized by Stueckelberg Stueckelberg (1941a, b, c), that if an electric charge is $-Q$ according to an observer with upward-pointing time orientation, it is $+Q$ according to an observer with downward-pointing time orientation (i.e. an anti-particle is a particle moving backward in time).
A closely analogous argument holds for the mass: if we use Gauss’s law to determine the mirror’s total (2-sided) Komar mass $\tilde{M}_{tot}$ by starting with the Killing one-form $k=-(1-\frac{2m}{r})dt$ (the one-form dual to the timelike Killing vector $\partial_{t}$), computing its “field strength” $\tilde{F}_{k}=dk$, and finding $\tilde{M}_{tot}\propto\int_{\partial\Sigma_{+}}\star\tilde{F}_{k}+\int_{\partial\Sigma_{-}}\star\tilde{F}_{k}$, the contributions from both sides of the mirror are now equal (rather than opposite, as they were for the black hole). In other words: in contrast to the Kruskal black hole, the black mirror’s Komar mass is positive in both exteriors, as expected from the positive-mass theorem Schoen and Yau (1979, 1981) – a much healthier state of affairs!
Now let us return to the surprising situation that, on the one hand, the black mirror’s total 2-sided electric charge $\tilde{Q}_{tot}$ (or total 2-sided mass $\tilde{M}_{tot}$), as determined by the total outward-pointing electric (or gravitational) flux, is non-zero; yet, on the other hand (and contrary to Gauss’s Law), there is no “enclosed” charge (or mass) confined to the black mirror (i.e. no delta-function stress energy tensor $T_{\mu\nu}$ or current density $j^{\mu}$ living on the mirror at $\sigma=0$), since the black mirror remains a smooth solution of the vacuum Einstein-Maxwell equations across $\sigma=0$ (and, all components of the Riemann curvature $R^{\alpha}_{\;\;\beta\gamma\delta}$, Ricci tensor $R_{\mu\nu}$, and Einstein tensor $G_{\mu\nu}$ remain smooth and finite across $\sigma=0$). Thus, the black mirror is a new incarnation of Misner and Wheeler’s idea of “mass without mass” and “charge without charge” and a new illustration of the previously-known fact that Stokes’ Theorem can be violated in certain topologically non-trivial circumstances Misner and Wheeler (1957); Sorkin (1977); Giulini (2018). In algebraic geometry, topologically nontrivial algebraic varieties like our black mirror are studied using standard “blow-up” techniques Griffiths and Harris (1978), and in follow-up work we are studying the black mirror from this standpoint to elucidate this interesting behavior.
Finally, consider a particle with electric charge $q$ following the trajectory $x^{\mu}(\tau)$ in the region $\sigma>0$ of the charged Kerr black mirror, and the image of this trajectory under the $CPT$ map $\varphi_{CPT}$, in the region $\sigma<0$. Given that the charge of the black mirror – as measured by an observer in $\sigma>0$ using Gauss’ law – is the same as the charge measured by an observer in $\sigma<0$, the invariance of the particle equations of motion under $CPT$ means that the charge of the particle as measured by an observer in $\sigma>0$ must also be the same as the charge of the CPT-image particle measured by an observer in $\sigma<0$, as shown in Figs. 4, 5. Similarly, the unusual topology in the stationary case (discussed in the preceding paragraphs) suggests that, although the two infalling “mirror particles” have he same charge, they nevertheless annihilate when they meet at the horizon. (This is also the picture obtained from smoothly extending the solution across the horizon, in Appendix C.) In follow-up work (mentioned above), we apply blow-up techniques to the black mirror to help clarify this fascinating situation.
Appendix C Geodesics
The geodesics of infalling particles in the black mirror spacetime are illustrated in Fig. 4. In this appendix we show that, working in the coordinates $(v_{+},\sigma,\theta,\varphi)$ that smoothly cross the black horizon, massive particle geodesics may be smoothly extended across the horizon to give mirror-symmetric trajectories like those illustrated in Fig. 4.
For the sake of brevity, throughout this appendix we drop the subscript on $v_{+}$ and adopt units where the Schwarzchild radius $2m$ is unity. The line element is then
$ds^{2}=-f(r)dv^{2}+2dvdr+r^{2}d\Omega^{2}$ (29)
with $f=1-1/r$. Since $k=\partial_{v}$ is a Killing vector, we have the conserved energy
$E=-k_{\mu}\dot{x}^{\mu}=f(r)\dot{v}-\dot{r},$ (30)
with $E$ positive ($E<1$ and $E>1$ for a bound and an unbound orbit, respectively). If, for simplicity, we restrict attention to radial geodesics and ignore angular motion, the mass-shell constraint $\dot{x}_{\mu}\dot{x}^{\mu}=-1$ for a massive particle becomes
$\dot{r}^{2}=E^{2}-f(r).$ (31)
Combining Eqs. (30, 31) we find, for an infalling particle,
$\frac{dv}{dr}=\frac{1-E/\sqrt{E^{2}-f(r)}}{f(r)},$ (32)
which may be integrated to yield
$v$ $=$ $v_{0}+r-{r^{1\over 2}E\sqrt{1+r(E^{2}-1)}\over E^{2}-1}$ (33)
$+$ ${E(3-2E^{2})\,{\rm sinh}^{-1}(r^{1\over 2}\sqrt{E^{2}-1})\over(E^{2}-1)^{3\over 2}}$ (34)
$+$ $2\ln(r^{1\over 2}E+\sqrt{1+r(E^{2}-1)}).$ (35)
where $v_{0}$ is an arbitrary constant. Now, if we re-express $r$ in terms $\sigma$ (in the units used in this appendix, $r=1+{1\over 4}\sigma^{2}$), we find that $v(\sigma)$ is a smooth, even function of $\sigma$, which “turns around” at horizon crossing. [^4] The explicit expression may prove useful in computing path integral propagators for quantum particles in a black mirror spacetime. This conclusion is unchanged when angular motion is included, for any infalling geodesic intersecting the black horizon.
Note that, for a point $P$ on the black horizon, the only points in its causal future are other (later) points on the horizon (and in particular, those on the null generator passing through $P$). So the correct physical interpretation of these extended (two-sided) geodesics is not that they describe particles passing from one exterior into the other; but rather, that (as shown in Figs. 4c and 4f) they describe a particle and its CPT mirror image meeting (and annihilating) at the black horizon. This, in turn, is reflected in the fact that, although the full (two-sided) geodesic is smoothly parameterized by $\sigma$, it is not smoothly parameterized by the proper time $\tau$ (since, as indicated by the arrows in Fig. 4), the proper time flows toward the black mirror on both sides.
Appendix D Saddle points of the action
@@
where dot denotes $\partial_{t}$ and prime $\partial_{x}$. Note that, we can derive this expression by simply extending the integrand (including $\sqrt{-g}$) smoothly from $x>0$ to $x<0$; or, equivalently, we can use the positive square root of $\sqrt{-g}$ if we also reverse the orientation of the $(\theta,\varphi)$ 2-sphere integration as we pass from $x>0$ to $x<0$; see Appendix B. Setting the variation with respect to $S$ zero yields the constraint that $\dot{L}=0$. (As will be seen later, in the solution $r^{\prime}/a$ is finite and non-vanishing everywhere.) Varying the action with respect to $L$ and then setting it to unity yields the constraint
@@
$ds^{2}=-a^{2}(x)\,dt^{2}+dx^{2}+r^{2}(x)d\Omega^{2},$ (43)
with $a=r^{\prime}$ and
${r^{\prime}}^{2}=1-\frac{2m}{r}+\frac{q_{e}^{2}}{r^{2}}-\Lambda r^{2}.$ (44)
The function $a=a(r)$ has at most three roots: $r_{-}$ which corresponds to the Cauchy horizon, $r_{+}$ which corresponds to the event horizon and $r_{c}$ which corresponds to the cosmological horizon, with $r_{c}=+\infty$ if $\Lambda\leq 0$. Since $a^{2}\geq 0$ it follows that the range of $r$ lies in the interval $[r_{+},r_{c}]$. Moreover, CPT symmetry implies that $r$ obeys $r^{\prime}(0)=0$, and therefore $r(0)$ is a root of $a=a(r)$. If $r(0)=r_{c}$ then the function $r=r(x)$ is a constant. Thus, the only physical initial conditions for $r$ are $r(0)=r_{+}$ and $r^{\prime}(0)=0$. Extending this solution to the horizon by changing coordinates from Schwarzschild time $t$ to Eddington-Finkelstein time $v$ ($v_{+}$ or $v_{-}$) again brings us back to the black mirror discussed in the main text.